How many mL of 0.378 M silver nitrate are needed to titrate all the chloride in 3.97g of pure [Co(en)2Cl2]
Answer #1

for titration moles of chlorine in [Co(en)2Cl2] = mole of AgNO3

so mole of [Co(en)2Cl2] = 3.97 /  250 = 0.01588 mol

now mole of Cl- = 2 * 0.01588 = 0.03176 mol

now since mole of AgNO3 = mole of Cl-

so volume = moles / concentation

= 0.03176 / 0.378 = 0.084 L = 84 ml